if x^3+px+r and 3x^2+p have a common factor, then p^3/27+r^2/4=0

Grade:8

1 Answers

Arun
5 years ago
 
Let x-a is the common factor.
Then putting X=a we get a^2= -p/3 from 2nd equn
Then 
Putting X=a in 1st eqn we get
a^3+pa+r=0
a(a^2+p)+r=0
a(-p/3+p)+r=0
a= - 3r/2p
 
Now put this value of a to the 1st eqn.
r^2/4 + p^3/27=0
 
Regards
Arun